OpenStudy (togundeji):

9. At the end of the season, a soccer team’s win-to-loss ratio is 5:3. If the team lost 21 games, how many games did they win? (1 point) 7 21 35 56 10. Mark and Hayden won a raffle and decide to split the prize money using the ratio 9:8. If Hayden received $64, which of the following was the total amount of money that Mark and Hayden won? (1 point) $72 $80 $136 $170

OpenStudy (togundeji):

@Aveline can you help?

OpenStudy (togundeji):

Are you good a mathematics, @legomyego180 ?

OpenStudy (anonymous):

\[\frac{ 5 }{ 3 }=\frac{ x }{ 21 }\] find x

OpenStudy (legomyego180):

^ Correct idea, these are proportions

OpenStudy (anonymous):

similarly 10 th 9+8=17 \[\frac{ 17 }{ 8 }=\frac{ x }{ 8 }\] find x

OpenStudy (togundeji):

So would the answer be 35 for question 9, and $80 for question 10? @surjithayer

OpenStudy (anonymous):

correction type error \[\frac{ 17 }{ 8 }=\frac{ x }{ 64 }\]

OpenStudy (togundeji):

ok

OpenStudy (togundeji):

This is for question 10?

OpenStudy (togundeji):

@surjithayer

OpenStudy (anonymous):

first we have to add ratios i.e.,9+8=17 yes

OpenStudy (togundeji):

oh okay

OpenStudy (anonymous):

answer for 9th is correct.

OpenStudy (togundeji):

Okay so that means the answer for question 10 is wrong

OpenStudy (togundeji):

Ok is question 10 136? @surjithayer

OpenStudy (anonymous):

correct.

OpenStudy (togundeji):

Thanks, mate! @surjithayer

OpenStudy (anonymous):

yw